# Characterizing graph-nonedge pairs with single interval Cayley configuration spaces in 3-dimensions

### v1

@@ C.3.3. Proof of Lemma 7 [Unpruned possible graphs for top-level $H^{-1}$]
Let $x$ be any vertex in $H^{-1}\setminus(C^{-1}\setminus\{u,v\})$ that neighbors $w_{2}$. Also, assume wlog that $w_{1}w_{2}$ is contracted in $[G\cup f]_{A}$. If every path in $G\cup f$ between either $u$ or $v$ and $w_{3}$ contains some vertex in $\{x,w_{1},w_{2}\}$, then the subgraph induced by these three vertices is an expanded $f$-separating cms of $G\cup f$ whose expanded $f$-component is a proper subgraph of $H^{-1}$, which is a contradiction. Otherwise, if some path in $G\cup f$ between either $u$ or $v$ and $w_{1}$ does not contain any vertex in $\{x,w_{2},w_{3}\}$, then an argument as in Case 1 of Lemma 39 shows that $xw_{2}$ is reducing, which is a contradiction. Else, note that $f$ is retained in the minor obtained by contracting $xw_{2}$, and so $xw_{2}$ is not of Type (3). Also, an argument as in Case 1 of Lemma 39 shows that $xw_{2}$ is not of Type (1). Hence, $xw_{2}$ of of Type (2). Combining our assumption on the paths in $G\cup f$ with an argument as in Case 1 of Lemma 39, we see that the subgraph of $G\cup f$ induced by $\{x,w_{2},w_{3}\}$ is an expanded $f$-separating cms of $G\cup f$ whose expanded $f$-component is a proper subgraph of $H^{-1}$, which is a contradiction.
@@
By our assumption and Lemma 5, any vertex in $H^{-1}\setminus(C^{-1}\setminus\{u,v\})$ neighbors some vertex in $\{w_{1},w_{3}\}$. Consider the case where some vertex $x$ in this subgraph neighbors exactly one vertex in $\{w_{1},w_{3}\}$, wlog say $w_{1}$. Also, assume that every path in $G\cup f$ between either $u$ or $v$ and $w_{3}$ contains some vertex in $\{x,w_{1},w_{2}\}$. If $w_{1}w_{2}$ is contracted in $[G\cup f]_{A}$, then the subgraph induced by these three vertices is an expanded $f$-separating cms of $G\cup f$ whose expanded $f$-component is a proper subgraph of $H^{-1}$, which is a contradiction. Otherwise, since $f$ is retained in the minor $[G\cup f]$ obtained by contracting $xw_{1}$, $xw_{1}$ is not of Type (3). Also, since $G\cup f$ is an atom it has some path between either $u$ or $v$ and $w_{2}$ that does not contain any vertex in $\{w_{1},w_{2}\}$. Hence, $[G\cup f]_{A}$ has an $[f]_{A}$-preserving forbidden minor, it is easy to see that $[G\cup f]$ has an $[f]$-preserving forbidden minor, and so $xw_{1}$ is not of Type (1). Consequently, $xw_{1}$ is of Type (2). We arrive at the same contradiction as above by using an argument as in Case 1 of Lemma 39.
@@
If either $H^{-1}\setminus(C^{-1}\setminus\{u,v\})$ contains at least three vertices or $u$ and $v$ each neighbor both $w_{1}$ and $w_{3}$, then $H^{-1}\cup w_{1}w_{3}$ has an $f$-preserving forbidden minor, which contradicts Lemma 36. Also, if exactly one vertex in $\{u,v\}$, wlog say $v$, neighbors either $w_{2}$ or exactly one vertex in $\{w_{1},w_{3}\}$, then an argument as in Case 1 shows that some edge incident on $v$ is reducing, which is a contradiction. Hence, assume none of these situations occur. Additionally, assume that $H^{-1}\setminus(C^{-1}\setminus\{u,v\})$ contains exactly two vertices, say $x$ and $y$. If all paths between either $u$ or $v$ and any vertex in $C^{-1}$ contains some vertex in $\{x,y\}$, then consider the minor $[G\cup f]$ obtained by contracting $vx$. The unique atom of $[G\cup f]$ that contains $[f]$ is $K_{3}$, and so $vx$ is not of Type (3). Also, since $[G\cup f]_{A}$ has an $[f]_{A}$-preserving forbidden minor, it is easy to see that $[G\cup f]$ has an $[f]$-preserving forbidden minor, and so $vx$ is not of Type (3). Consequently, $vx$ is of Type (2). Using as argument as in Case 1 of Lemma 39, we see that the subgraph of $G\cup f$ induced by $\{v,x,y\}$ is an expanded $f$-separating cms of $G\cup f$ whose expanded $f$-component is a proper subgraph of $H^{-1}$, which is a contradiction.
@@
Let $C^{-1}$ be a top-level expanded $f$-separating cms of a minimal graph-nonedge pair $(G,f)$, and let its expanded $f$-component be $H^{-1}$. If $C^{-1}$ is the graph in Figure 13(b), 13(c), or 13(d), then $H^{-1}$ is one of the graphs in Figure 15 not mentioned in Lemma 39 such that $f$ is the red edge and $C^{-1}$ is the subgraph induced by $\{w_{i}\}$.
The proof of Lemma 41 is omitted here since it is very similar to the proofs of Lemmas 39 and 40, but it is available upon request.
@@
The lemma follows immediately from Lemmas 6, 39, 40, and 41. ∎
C.3.4. Proof of Proposition 8 [Minimal pair inductive structure]
Consider a minimal graph-nonedge pair $(G,f)$ with an expanded $f$-separating cms $C^{-1}$ whose expanded $f$-component is $H^{-1}$. First, we prove Lemma 44, below, which eliminates some of the graphs in Figure 15 as possibilities for $H^{-1}$. The proof of this lemma requires Lemmas 42 and 43, also below.
Figure 20. Graphs in the proof of Lemma 42.
@@ C.3.4. Proof of Proposition 8 [Minimal pair inductive structure]
By the definition of $C^{-1}$, $C$ is an $f$-separating cms of some minor $[G\cup f]$, and the minor component $I$ of $C$ has an $[f]$-preserving forbidden minor $[I]_{A}$. Note that each nonedge $f^{\prime}$ of $C^{-1}$ is preserved in $[G\cup f]$. Hence, it suffices to show that $I$ has an $[f^{\prime}]$-preserving forbidden minor, which we demonstrate now. By Lemma 6, $C^{-1}$ is one of the graphs in Figures 13(a)-13(d). This implies that $C$ is either $K_{2}$ or $K_{3}$. If $C$ is $K_{2}$, then $[I]_{A}$ is $[f^{\prime}]$-preserving, since it is $[f]$-preserving. Otherwise, let $[I]_{B}$ be the minor obtained by contracting all edges that are contracted in $[I]_{A}$ except for any edge of $C$. If $[I]_{B}=[I]_{A}$, then we are done. Otherwise, since $[I]_{A}$ is $[f]$-preserving, it is obtained from $[I]_{B}$ by contracting exactly one edge of $[C]_{B}$. Since $[I]_{A}$ is either $K_{5}$ or $K_{2,2,2}$, this implies that some spanning subgraph of $[I]_{B}$ is one of the graphs in Figure 20 such that $[C]_{B}$ is the subgraph induced by $\{x_{i}\}$.
Next, assume wlog that contracting $x_{1}x_{2}$ yields $[I]_{A}$. If the above-mentioned some spanning subgraph of $[I]_{B}$ is the graph in Figure 20(a), then $[I]_{A}$ is $K_{5}$. This implies that $x_{3}$ neighbors each vertex in $\{y_{1},y_{2},y_{3}\}$, the union of the neighborhoods of $x_{1}$ and $x_{2}$ is $\{y_{1},y_{2},y_{3},x_{3}\}$, and wlog $x_{1}$ has at least two neighbors in $\{y_{1},y_{2},y_{3}\}$, say $y_{1}$ and $y_{2}$. For all cases of the neighborhoods of $x_{1}$ and $x_{2}$, we can clearly obtain an $[f^{\prime}]_{B}$-preserving $K_{5}$ minor of $[I]_{B}$, and hence an $[f^{\prime}]$-preserving $K_{5}$ minor of $I$.
Finally, if the above-mentioned some spanning subgraph of $[I]_{B}$ is the graph in Figure 20(b), then $[I]_{A}$ is $K_{2,2,2}$. This implies wlog that $x_{3}$ neighbors each vertex in $\{y_{1},y_{3},y_{4}\}$, the union of the neighborhoods of $x_{1}$ and $x_{2}$ is $\{y_{1},y_{2},y_{3},x_{3}\}$, and $x_{1}$ wlog has at least two neighbors in $\{y_{1},y_{2},y_{3}\}$, say $y_{1}$ and $y_{2}$. For all cases of the neighborhoods of $x_{1}$ and $x_{2}$, we can clearly obtain an $[f^{\prime}]_{B}$-preserving forbidden minor of $[I]_{B}$, and hence an $[f^{\prime}]$-preserving forbidden minor of $I$. This completes the proof. ∎
@@
Lemma 42 states that the expanded minor component of $C^{-1}$ has an $f^{\prime}$-preserving forbidden minor for each nonedge $f^{\prime}$ of $C^{-1}$. Hence, Lemma 27 shows that some atom $J$ of $G^{\prime}\cup f^{\prime}$ has an $f^{\prime}$-preserving forbidden minor. First, we demonstrate that we can choose $f^{\prime}$ to be contained in $J$. Note that $G\cup f$ is an atom, $(G\cup f)\setminus C^{-1}$ is disconnected, and $G^{\prime}$ is the union of $C^{-1}$ and some non-empty set of connected components of $(G\cup f)\setminus C^{-1}$. Furthermore, by inspecting all possibilities for $C^{-1}$ given by Lemma 6, we see that either $C^{-1}$ is an atom or, for any cms $E$ of $C^{-1}$, there are exactly two $E$-components of $C^{-1}$, both of which are atoms. Therefore, Lemma 63 in Appendix E shows that the atom graph of $G^{\prime}$ is a path.
Next, if $G^{\prime}$ is an atom, then Lemma 59 Statement (1) in Appendix E shows that any choice of $f^{\prime}$ is contained in $J$. Note that if $C^{-1}$ is the graph in Figure 13(b), then $C^{-1}$ is an atom. Hence, Lemma 61 in Appendix E shows that $G^{\prime}$ is an atom. Combining these facts demonstrates that any choice for $f^{\prime}$ is contained in $J$. Otherwise, if $G^{\prime}$ is not an atom, then the atom graph of $G^{\prime}$ has two distinct leaves $y_{1}$ and $y_{2}$ with not necessarily distinct neighbors $x_{1}$ and $x_{2}$, respectively. Using an argument similar to the proof of Lemma 63, we see that $G^{\prime}(y_{1})\setminus G^{\prime}(x_{1})$ contains some vertex $u$ of $C^{-1}$, $G^{\prime}(y_{2})\setminus G^{\prime}(x_{2})$ contains some vertex $v$ of $C^{-1}$, and $G^{\prime}(y_{1})\setminus G^{\prime}(x_{1})$ and $G^{\prime}(y_{2})\setminus G^{\prime}(x_{2})$ do not share any vertices. These facts show that $uv$ is a nonedge of $C^{-1}$. Therefore, Lemma 59 Statement (2) shows that $J$ contains $uv$.
@@
Lemma 44 (Pruning cases for $H^{-1}$).
The expanded $f$-component $H^{-1}$ of any expanded $f$-separating cms $C^{-1}$ of a minimal graph-nonedge pair is one of the graphs in Figures 15(a) - 15(m), where $C^{-1}$ is the subgraph induced by the vertex set $\{w_{i}\}$. Furthermore, if $H^{-1}$ is one of the graphs in Figures 15(c), 15(d), 15(e), 15(i), 15(k), or 15(l), then at least one dashed line-segment is a nonedge.
Proof.
By Proposition 7, $H^{-1}$ is one of the graphs in Figure 15. Let $G^{\prime}=G\setminus(H^{-1}\setminus C^{-1})$ and consider the nonedge $f^{\prime}$ of $C^{-1}$ and the atom $J$ of $G^{\prime}\cup f^{\prime}$ given by Lemma 43. For each case of $H^{-1}$ in Figures 15(n) - 15(r) or Figures 15(c), 15(d), 15(i), 15(k), or 15(l) such that every dashed line-segment is an edge, observe that there exists a minor $[G\cup f]$ obtained via a single contraction such that $f$ is retained in $[G\cup f]$, $[f]=[f^{\prime}]$, and some induced subgraph of $[G\cup f]$ is $G^{\prime}\cup f^{\prime}$ such that $[f]$ is $f^{\prime}$. These facts along with the existence of $J$ show that the edge contracted in $[G\cup f]$ is reducing with respect to $(G,f)$, which contradicts the minimality of this pair. ∎
Next, we prove Lemma 47, below, using Lemmas 45 and 46, also below.
Lemma 45 (Properties of graphs with $H^{-1}$ in Figures 15(a) - 15(m)).
Let $G$ be a graph with subgraphs $G_{1}$, $G_{2}$, and $S$ such that $G=G_{1}\cup G_{2}$, $G_{1}$ is one of the graphs in Figures 15(a) - 15(m), $S=G_{1}\cap G_{2}$ is the subgraph of $G_{1}$ induced by some subset of $\{w_{i}\}$ that contains some nonedge $e$, and $G\setminus S$ is disconnected. If $G_{2}\cup e$ is an atom and no clique subgraph of $S$ separates the endpoints of $e$ in $G_{2}$, then $G$ is an atom.

### v2

@@ B.2.1   Proof of Lemma 31 [Unpruned possible graphs for top-level $H^{-1}$]
Let $x$ be any vertex in $H^{-1}\setminus(C^{-1}\setminus\{u,v\})$ that neighbors $w_{2}$. Also, assume wlog that $w_{1}w_{2}$ is contracted in $[G\cup f]_{1}$. If every path in $G\cup f$ between either $u$ or $v$ and $w_{3}$ contains some vertex in $\{x,w_{1},w_{2}\}$, then the subgraph induced by these three vertices is an expanded $f$-separating CMS of $G\cup f$ whose expanded $f$-component is a proper subgraph of $H^{-1}$, which is a contradiction. Otherwise, if some path in $G\cup f$ between either $u$ or $v$ and $w_{1}$ does not contain any vertex in $\{x,w_{2},w_{3}\}$, then an argument as in Case 1 of Lemma 49 shows that $xw_{2}$ is reducing, which is a contradiction. Else, note that $f$ is retained in the minor obtained by contracting $xw_{2}$, and so $xw_{2}$ is not of Type (3). Also, an argument as in Case 1 of Lemma 49 shows that $xw_{2}$ is not of Type (1). Hence, $xw_{2}$ of of Type (2). Combining our assumption on the paths in $G\cup f$ with an argument as in Case 1 of Lemma 49, we see that the subgraph of $G\cup f$ induced by $\{x,w_{2},w_{3}\}$ is an expanded $f$-separating CMS of $G\cup f$ whose expanded $f$-component is a proper subgraph of $H^{-1}$, which is a contradiction.
@@
By our assumption and Lemma 28, any vertex in $H^{-1}\setminus(C^{-1}\setminus\{u,v\})$ neighbors some vertex in $\{w_{1},w_{3}\}$. Consider the case where some vertex $x$ in this subgraph neighbors exactly one vertex in $\{w_{1},w_{3}\}$, wlog say $w_{1}$. Also, assume that every path in $G\cup f$ between either $u$ or $v$ and $w_{3}$ contains some vertex in $\{x,w_{1},w_{2}\}$. If $w_{1}w_{2}$ is contracted in $[G\cup f]_{1}$, then the subgraph induced by these three vertices is an expanded $f$-separating CMS of $G\cup f$ whose expanded $f$-component is a proper subgraph of $H^{-1}$, which is a contradiction. Otherwise, since $f$ is retained in the minor $[G\cup f]$ obtained by contracting $xw_{1}$, $xw_{1}$ is not of Type (3). Also, since $G\cup f$ is an atom it has some path between either $u$ or $v$ and $w_{2}$ that does not contain any vertex in $\{w_{1},w_{2}\}$. Hence, $[G\cup f]_{1}$ has an $[f]_{1}$-preserving forbidden minor, it is easy to see that $[G\cup f]$ has an $[f]$-preserving forbidden minor, and so $xw_{1}$ is not of Type (1). Consequently, $xw_{1}$ is of Type (2). We arrive at the same contradiction as above by using an argument as in Case 1 of Lemma 49.
@@
If either $H^{-1}\setminus(C^{-1}\setminus\{u,v\})$ contains at least three vertices or $u$ and $v$ each neighbor both $w_{1}$ and $w_{3}$, then $H^{-1}\cup w_{1}w_{3}$ has an $f$-preserving forbidden minor, which contradicts Lemma 46. Also, if exactly one vertex in $\{u,v\}$, wlog say $v$, neighbors either $w_{2}$ or exactly one vertex in $\{w_{1},w_{3}\}$, then an argument as in Case 1 shows that some edge incident on $v$ is reducing, which is a contradiction. Hence, assume none of these situations occur. Additionally, assume that $H^{-1}\setminus(C^{-1}\setminus\{u,v\})$ contains exactly two vertices, say $x$ and $y$. If all paths between either $u$ or $v$ and any vertex in $C^{-1}$ contains some vertex in $\{x,y\}$, then consider the minor $[G\cup f]$ obtained by contracting $vx$. The unique atom of $[G\cup f]$ that contains $[f]$ is $K_{3}$, and so $vx$ is not of Type (3). Also, since $[G\cup f]_{1}$ has an $[f]_{1}$-preserving forbidden minor, it is easy to see that $[G\cup f]$ has an $[f]$-preserving forbidden minor, and so $vx$ is not of Type (3). Consequently, $vx$ is of Type (2). Using as argument as in Case 1 of Lemma 49, we see that the subgraph of $G\cup f$ induced by $\{v,x,y\}$ is an expanded $f$-separating CMS of $G\cup f$ whose expanded $f$-component is a proper subgraph of $H^{-1}$, which is a contradiction.
@@
Let $C^{-1}$ be a top-level expanded $f$-separating CMS of a minimal graph-nonedge pair $(G,f)$, and let its expanded $f$-component be $H^{-1}$. If $C^{-1}$ is the graph in Figure 22(b), 22(c), or 22(d), then $H^{-1}$ is one of the graphs in Figure 20 not mentioned in Lemma 49 with $f$ as the red edge and $C^{-1}$ as the subgraph induced by $\{w_{i}\}$.
The proof of Lemma 51 is omitted here since it is very similar to the proofs of Lemmas 49 and 50, but it is available upon request.
@@
The lemma follows immediately from Lemmas 30 and 49-51. ∎
Lemma 32 in Section 5.1 is proved using Lemma 53, below. The proof of Lemma 53 requires Lemma 52, also below.
Figure 26: Graphs in the proof of Lemma 52.
@@
By the definition of $C^{-1}$, $C$ is an $f$-separating CMS of some minor $[G\cup f]$, and the minor component $I$ of $C$ has an $[f]$-preserving forbidden minor $[I]_{1}$. Note that each nonedge $f^{\prime}$ of $C^{-1}$ is preserved in $[G\cup f]$. Hence, it suffices to show that $I$ has an $[f^{\prime}]$-preserving forbidden minor, which we demonstrate now. By Lemma 30, $C^{-1}$ is one of the graphs in Figures 22(a)-22(d). This implies that $C$ is either $K_{2}$ or $K_{3}$. If $C$ is $K_{2}$, then $[I]_{1}$ is $[f^{\prime}]$-preserving, since it is $[f]$-preserving. Otherwise, let $[I]_{2}$ be the minor obtained by contracting all edges that are contracted in $[I]_{1}$ except for any edge of $C$. If $[I]_{2}=[I]_{1}$, then we are done. Otherwise, since $[I]_{1}$ is $[f]$-preserving, it is obtained from $[I]_{2}$ by contracting exactly one edge of $[C]_{2}$. Since $[I]_{1}$ is either $K_{5}$ or $K_{2,2,2}$, this implies that some spanning subgraph of $[I]_{2}$ is one of the graphs in Figure 26 with $[C]_{2}$ as the subgraph induced by $\{x_{i}\}$.
Next, assume wlog that contracting $x_{1}x_{2}$ yields $[I]_{1}$. If the above-mentioned some spanning subgraph of $[I]_{2}$ is the graph in Figure 26(a), then $[I]_{1}$ is $K_{5}$. This implies that $x_{3}$ neighbors each vertex in $\{y_{1},y_{2},y_{3}\}$, the union of the neighborhoods of $x_{1}$ and $x_{2}$ is $\{y_{1},y_{2},y_{3},x_{3}\}$, and wlog $x_{1}$ has at least two neighbors in $\{y_{1},y_{2},y_{3}\}$, say $y_{1}$ and $y_{2}$. For all cases of the neighborhoods of $x_{1}$ and $x_{2}$, we can clearly obtain an $[f^{\prime}]_{2}$-preserving $K_{5}$ minor of $[I]_{2}$, and hence an $[f^{\prime}]$-preserving $K_{5}$ minor of $I$.
Finally, if the above-mentioned some spanning subgraph of $[I]_{2}$ is the graph in Figure 26(b), then $[I]_{1}$ is $K_{2,2,2}$. This implies wlog that $x_{3}$ neighbors each vertex in $\{y_{1},y_{3},y_{4}\}$, the union of the neighborhoods of $x_{1}$ and $x_{2}$ is $\{y_{1},y_{2},y_{3},x_{3}\}$, and $x_{1}$ wlog has at least two neighbors in $\{y_{1},y_{2},y_{3}\}$, say $y_{1}$ and $y_{2}$. For all cases of the neighborhoods of $x_{1}$ and $x_{2}$, we can clearly obtain an $[f^{\prime}]_{2}$-preserving forbidden minor of $[I]_{2}$, and hence an $[f^{\prime}]$-preserving forbidden minor of $I$. This completes the proof. ∎
@@
Lemma 52 states that the expanded minor component of $C^{-1}$ has an $f^{\prime}$-preserving forbidden minor for each nonedge $f^{\prime}$ of $C^{-1}$. Hence, Lemma 22 shows that some atom $J$ of $G^{\prime}\cup f^{\prime}$ has an $f^{\prime}$-preserving forbidden minor. First, we demonstrate that we can choose $f^{\prime}$ to be contained in $J$. Note that $G\cup f$ is an atom, $(G\cup f)\setminus C^{-1}$ is disconnected, and $G^{\prime}$ is the union of $C^{-1}$ and some non-empty set of connected components of $(G\cup f)\setminus C^{-1}$. Furthermore, by inspecting all possibilities for $C^{-1}$ given by Lemma 30, we see that either $C^{-1}$ is an atom or, for any CMS $E$ of $C^{-1}$, there are exactly two $E$-components of $C^{-1}$, both of which are atoms. Therefore, Lemma 65 in C shows that the atom graph of $G^{\prime}$ is a path.
Next, if $G^{\prime}$ is an atom, then Lemma 61 Statement (1) in C shows that any choice of $f^{\prime}$ is contained in $J$. Note that if $C^{-1}$ is the graph in Figure 22(b), then $C^{-1}$ is an atom. Hence, Lemma 63 in C shows that $G^{\prime}$ is an atom. Combining these facts demonstrates that any choice for $f^{\prime}$ is contained in $J$. Otherwise, if $G^{\prime}$ is not an atom, then the atom graph of $G^{\prime}$ has two distinct leaves $y_{1}$ and $y_{2}$ with not necessarily distinct neighbors $x_{1}$ and $x_{2}$, respectively. Using an argument similar to the proof of Lemma 65, we see that $G^{\prime}(y_{1})\setminus G^{\prime}(x_{1})$ contains some vertex $u$ of $C^{-1}$, $G^{\prime}(y_{2})\setminus G^{\prime}(x_{2})$ contains some vertex $v$ of $C^{-1}$, and $G^{\prime}(y_{1})\setminus G^{\prime}(x_{1})$ and $G^{\prime}(y_{2})\setminus G^{\prime}(x_{2})$ do not share any vertices. These facts show that $uv$ is a nonedge of $C^{-1}$. Therefore, Lemma 61 Statement (2) shows that $J$ contains $uv$.
@@
Proof of Lemma 32.
By Lemma 31, $H^{-1}$ is one of the graphs in Figure 20. Let $G^{\prime}=G\setminus(H^{-1}\setminus C^{-1})$ and consider the nonedge $f^{\prime}$ of $C^{-1}$ and the atom $J$ of $G^{\prime}\cup f^{\prime}$ given by Lemma 53. For each case of $H^{-1}$ in Figures 20(n) - 20(r) or Figures 20(c), 20(d), 20(i), 20(k), or 20(l) such that every dashed line-segment is an edge, observe that there exists a minor $[G\cup f]$ obtained via a single contraction such that $f$ is retained in $[G\cup f]$, $[f]=[f^{\prime}]$, and some induced subgraph of $[G\cup f]$ is $G^{\prime}\cup f^{\prime}$ with $[f]$ as $f^{\prime}$. These facts along with the existence of $J$ show that the edge contracted in $[G\cup f]$ is reducing with respect to $(G,f)$, which contradicts the minimality of this pair. ∎
The next subsection is dedicated to proving Lemma 33 from Section 5.1.
B.2.2   Proof of Lemma 33
The proof requires Lemmas 54 and 55, below.
Lemma 54 (Properties of graphs with $H^{-1}$ in Figures 20(a) - 20(m)).
Let $G$ be a graph with subgraphs $G_{1}$, $G_{2}$, and $S$ such that $G=G_{1}\cup G_{2}$, $G_{1}$ is one of the graphs in Figures 20(a) - 20(m), $S=G_{1}\cap G_{2}$ is the subgraph of $G_{1}$ induced by some subset of $\{w_{i}\}$ that contains some nonedge $e$, and $G\setminus S$ is disconnected. If $G_{2}\cup e$ is an atom and no clique subgraph of $S$ separates the endpoints of $e$ in $G_{2}$, then $G$ is an atom.